OpenStudy (anonymous):

Last, please? Thank you. :) Two-column proof. Given: BA⊥BD DF⊥BD BC||DE Proof: ∠1 is complementary to ∠3.

OpenStudy (anonymous):



OpenStudy (anonymous):

first,the question is incorrect,BF⊥BD is incorrect,DF⊥BD is correct, wait a moment,

OpenStudy (anonymous):

I will prove it :)

OpenStudy (anonymous):

PFEEH is right !

OpenStudy (anonymous):

wrong? Why?

OpenStudy (anonymous):

What is the BF ?!!!

OpenStudy (anonymous):

wait let me check

OpenStudy (anonymous):

BF⊥BD?????think about it! It's not triangle...

OpenStudy (anonymous):

Oh wait! You're right. It should be DF.

OpenStudy (anonymous):

PFEH : you re right !

OpenStudy (anonymous):

Let me change it

OpenStudy (anonymous):

Please draw a new picture !

OpenStudy (anonymous):

found It!!!

OpenStudy (anonymous):

\[D _{3} = B _{2}\] ok?

OpenStudy (anonymous):

@stupidinmath Ok?

OpenStudy (anonymous):



OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

No @E.ali ! Don't tell him the answer in that case!

OpenStudy (anonymous):

No it wasnt answer !

OpenStudy (anonymous):

Its lright. I have to do the table proof anyways

OpenStudy (anonymous):

then, \[90-B _{2} = 90 - D _{4}\]

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

Yep, got it:)

OpenStudy (anonymous):

so, \[-B _{2} = -D _{4} -> B _{2} = D _{4}\]

OpenStudy (anonymous):

we give AC || EF || BD In ABC we have : A=90 B+C=90 A+B+C=90 B+C = B2

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

@E.ali !!!!! Don't tell himthe answer!!!

OpenStudy (anonymous):

Oh.. okay..

OpenStudy (anonymous):

It the proof ! He should know that ! He can answer than me ! Look !>>

OpenStudy (anonymous):

Oh Oh,sorry @stupidinmath , i made a mistake!

OpenStudy (anonymous):

I think so, too. And, its a "SHE"

OpenStudy (anonymous):

Its alright:)

OpenStudy (anonymous):

\[D _{4} = B _{1} \]

OpenStudy (anonymous):

∠1 is complementary to ∠2 ∠3 is complementary to ∠4 so we have: ∠1 is complementary to ∠3

OpenStudy (anonymous):

OK? got it?

OpenStudy (anonymous):

Yes,very clearly